# Chance Constrained PDE-Constrained Optimal Design Strategies Under High-Dimensional Uncertainty

### v1

@@ 2.1 Forward thermomechanical model and design objectives
$r(\mathbf{u},\mathbf{v},m,d):=\,_{\mathcal{V}}\langle\mathbf{v},\mathcal{R}(\mathbf{u},m,d)\rangle_{\mathcal{V}^{\prime}}=0\quad\forall\mathbf{v}\in\mathcal{V}$ (7)
where ,$\,{}_{\mathcal{V}}\langle\cdot,\cdot\rangle_{\mathcal{V}^{\prime}}$ represent the duality pairing between the $\mathcal{V}$ and $\mathcal{V}^{\prime}$.
Figure 2: Finite element solution of the forward model for the beam-insulator system with constant porosity $\phi_{f}=0.7$, depicting the solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$, displacement magnitude $||\mathbf{u}_{s}||$, and von Mises stress $T_{\text{VM}}$.
Figure 2 presents the deterministic finite element solution of the forward model for the beam-insulator system under constant porosity $\phi_{f}=0.7$, utilizing the model parameters listed in Table 1. Thermal expansion of the thermoelastic concrete beam induces stress concentrations within the aerogel thermal break, risking mechanical failure. This can be mitigated by varying the aerogel’s porosity across the insulation, maintaining its effectiveness in interrupting heat transfer between the building interior and exterior.
Design objectives
The design objective $Q$ is defined as thermal compliance of the aerogel thermal break component over domain $\Omega$,
$Q$ $=$ $\frac{1}{2}\sum_{i=s,f}\big{<}\phi_{i}\,\kappa_{i}\,\nabla\theta_{i},\nabla\theta_{i}\big{>}+{\big{<}\phi_{i}h_{\mathrm{air}}(\theta_{i}-\theta_{\mathrm{amb}}),\theta_{i}\big{>}},$
The design problem focuses on determining the spatial distribution of the fluid volume fraction (porosity), $\phi_{f}(\mathbf{x})$, to minimize stress concentrations within the thermal break, thereby preventing mechanical failure while preserving its thermal insulation performance.
2.2 Uncertain design parameter
Due to the inherent challenges in precisely controlling the properties of additively manufactured aerogel, considerable uncertainty arises in the porosity values within the thermal break. To effectively model uncertainty, a design parameter as a spatial field $d\in[0,1]$ is introduced, employing a sigmoid mapping function, ensuring that porosity values asymptotically remain within the physically valid range, maintaining near-linear behavior within $\phi_{f}\in[90\%,10\%]$, and minimize the impact on the spatial correlation length, such that [31]
$\phi_{f}\equiv\text{sigmoid}\left(d+m\right).$ (9)
The uncertain parameter $m$ is represented as a Gaussian random field characterized by the Matérn covariance kernel. Specifically, $m$ is distributed as $\mathcal{N}(\bar{m},\mathcal{C})$ where $\bar{m}$ is the mean and $\mathcal{C}$ is the covariance operator. The sampling of Gaussian random field $m$ is equivalent to solving the elliptic equation with a robin boundary condition,
@@ 2.2 Uncertain design parameter
where $\nabla$ is gradient operator,$\Delta$ is the Laplace operator and $\mathbf{n}$ is the outward unit normal vector. Here, $W$ represents the spatial White Gaussian noise with a unit variance, and the parameters $\gamma$ and $\delta$ control the variance $\sigma^{2}={1}/{4\pi\,\delta\,\gamma}$ and spatial correlation length $L_{c}=\sqrt{{8\gamma}/{\delta}}$. The factor ${\sqrt{\delta\,\gamma}}/{1.42}$ (Robin coefficient) is chosen to minimize boundary artifacts [46], and $\boldsymbol{\Theta}$ is a symmetric positive definite two-dimensional (2D) anisotropic tensor
[…]
@@ 2.3 Chance-constrained PDE-constrained optimal design under uncertainty
We consider the design problem as minimizing the design objective, thermal compliance in (LABEL:eq:qois), while accounting for uncertainty in objective through its explicit dependence on the uncertain parameter $m$, i.e., $Q(d,m)$. Extending the mean-variance optimization, e.g., [18, 19, 31], the cost functional for the proposed optimal design takes the form,
$\mathcal{J}(d)=\mathbb{E}[Q(d,m)]+\beta_{V}\,\mathbb{V}[Q(d,m)]+\beta_{R}\,R(d).$ (12)
To prevent mechanical failure of the aerogel thermal break due to stress concentration, the maximum von Mises stress $T_{\text{VM}}$ must remain below a critical threshold. This requirement is incorporated as a probabilistic (chance) constraint in the optimization problem, as $T_{\text{VM}}$ depends on uncertain parameters. However, due to the non-differentiable nature of the maximum stress, direct optimization poses significant challenges. This challenge is addressed using stress aggregation function based on p-norm [41], which is a smooth approximation of the maximum stress and enables gradient-based optimization techniques.The p-norm for the von Mises stress can be computed as
@@
$f=T_{cr}-T_{\text{pn}},$ (14)
where $T_{cr}$ is the limiting critical stress and $T_{\text{pn}}$ is the p-norm of the von Mises stress over the domain. To avoid stress concentration, we consider a chance constraint:
$P(f(m,d)\geq 0)\leq\alpha_{c},$ (15)
for a critical chance $10<\alpha_{c}<$ 1. The probability is given by
$P(f(m,d)\geq 0)$ $=$ $\mathbb{E}[\mathbb{I}_{[0,\infty]}(f(m,d))]$ (16)
[…]
where $\mathbb{I}_{[0,\infty]}(f(m,d))$ is an indicator function defined as:
$\mathbb{I}_{[0,\infty]}(f(m,d))=\begin{cases}1&\text{if }f(m,d)\geq 0\\ 0&\text{if }f(m,d)<0\\ \end{cases}$ (17)
uncertainty as,
$d_{\text{opt}}(\mathbf{x})=\min_{{d}}$ $\ \mathcal{J}({d})$ (18)
@@
$P(f(m,d)\geq 0)\leq\alpha_{c}\quad\mathrm{in\ }\Omega$
Monte Carlo estimation
A common approach to compute the mean and variance terms in (12) is the Monte Carlo estimation for $N_{mc}$ samples,
@@
where $\{m^{(i)}\}_{i=1}^{N_{mc}}$ are independent and identically distributed (i.i.d) random samples from the uncertain and spatially correlated parameter. Similarly, the chance constraint function can be approximated by
$P(f(.,d)\geq 0):=\frac{1}{M_{f}}\sum_{i=1}^{M_{f}}\mathbb{I}_{[0,\infty)]}(f(m_{i},d))$ (21)
To achieve an accurate approximation, particularly for $f_{M}$ for $\alpha$ close to 00, a large number of samples are required due to slow convergence of the Monte Carlo method. The convergence rate of Monte Carlo method is $\mathcal{O}\big{(}\frac{1}{\sqrt{N_{mc}}}\big{)}$, therefore for accurate computation of the cost functional (12), a considerable number of PDE model evaluations will be required, making the optimization under uncertainty problem computationally prohibitive.
3 Scalable framework for chance constrained optimal design
In this section, we present a computationally tractable approach by approximating the mean and variance of the design objective by means of low-order Taylor expansion [18, 19]. The quadratic Taylor approximation of $Q$ at the mean of the uncertain parameter $\bar{m}$, assuming $Q$ is sufficiently smooth and Fréchet differentiable with respect to uncertain parameter is derived as,
$Q_{\text{QUAD}}(m)$ $=$ $\sum_{l=0}^{2}\partial^{l}_{m}Q(\bar{m})\,(m-\bar{m})^{l}$ (22)
$=$ $\bar{Q}+\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{Q}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}.$
where $\partial^{l}_{m}Q(\bar{m})$ is the $l$-th order tensor derivative at $\bar{m}$ and the objective $Q$ and its gradient and Hessian with respect to $m$, evaluated at mean $\bar{m}$, denoted as $\bar{Q}$, $\bar{Q}^{m}$ and $\bar{Q}^{mm}$, respectively. The close form for the expectation of the quadratic Taylor approximation in (22), considering $\mathbb{E}[\bar{Q}]=\bar{Q}$ and $\mathbb{E}[\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}]=0$, and Lemma 1 of [47], is obtained as
$\mathbb{E}[Q_{\text{QUAD}}]=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)},$ (23)
where ¯$\bar{\mathcal{H}}_{c}=\mathcal{C}\,\bar{Q}^{mm}$ denotes the covariance preconditioned Hessian computed at the mean of $m$ and $\mathrm{tr}(\cdot)$ is trace operator. Moreover, noting that variance is invariant with respect to its deterministic translation, e.g., $\mathbb{V}(m-\mathrm{const.})=\mathbb{V}(m)$), one can write,
$\mathbb{E}[(Q_{\text{QUAD}}(m)-\bar{Q})^{2}]=\mathbb{E}[(\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{Q}^{mm}(m-\bar{m}),\,m-\bar{m}\big{>})^{2}],$ (24)
which, upon expansion, allows each term to be expressed as
$\mathbb{E}[(\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>})^{2}]$ $=$ $\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>},$
$\mathbb{E}[(\big{<}\bar{Q}^{mm}(m-\bar{m}),\,m-\bar{m}\big{>})^{2}]$ $=$ $\mathrm{tr}(\bar{\mathcal{H}}_{c})^{2}+2\,\mathrm{tr}((\bar{\mathcal{H}}_{c})^{2}),$
$\mathbb{E}[\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}\,\big{<}\bar{Q}^{mm}(m-\bar{m}),\,m-\bar{m}\big{>}]$ $=$ $0$ (25)
Furthermore, the square of the first moment of $Q_{\text{QUAD}}(m)-\bar{Q}$ can be written as,
$\big{(}\mathbb{E}[Q_{\text{QUAD}}(m)-\bar{Q}]\big{)}^{2}=\big{(}\mathbb{E}[Q_{\text{QUAD}}(m)]-\mathbb{E}[\bar{Q}]\big{)}^{2}=\frac{1}{4}\big{(}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)}\big{)}^{2}$ (26)
From (25) and (26), the variance of the quadratic Taylor approximation of $Q$ in (22) is obtained as,
$\mathbb{V}[Q_{\text{QUAD}}(m)]=\mathbb{V}[Q_{\text{QUAD}}(m)-\bar{Q}]=\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{c})^{2}).$ (27)
Consequently, the mean and variance of the design objective in (12) can be replaced by their corresponding quadratic Taylor approximations in (23) and (27), respectively, leading to the approximated design objective as
$\mathcal{J}_{\mathrm{QUAD}}(d)=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)}+\beta_{V}\,\bigg{(}\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{c})^{2})\bigg{)}+\beta_{R}\,R(d).$ (28)
@@ 3 Scalable framework for chance constrained optimal design
Similar to (22), we can construct a second-order Taylor expansion of the constraint function $f$ at $\bar{m}$ as,
$f_{QUAD}(m)=\bar{f}+\big{<}\bar{f}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{f}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}.$ (29)
@@
The probability $P(f(.,d)\geq 0)$ by Monte Carlo estimator (21) with the quadratic approximation yields,
$P(f(m,d))\geq 0)=\frac{1}{M_{f}}\mathbb{I}_{[0,\infty)}(f_{\text{QUAD}}(m_{i},d)).$ (30)
This approximation of the constraint function obviates the necessity for a large number of PDE solves to assess the probability utilizing sample averaging methods. In upcoming sections, we present the Lagrangian formalism to derive the $m$-gradient and $m$-Hessian, representing the gradient and Hessian of $Q$ with respect to $m$. We also introduce an efficient method for evaluating the covariance preconditioned Hessian, ¯$\bar{\mathcal{H}}_{c}$.
3.1 Analytical gradient and hessian of design objective with respect to the uncertain parameter
The Lagrangian functional, considering the adjoint variable $\mathbf{v}^{Q}$ as the Lagrange multiplier associated with the state equation with respect to the objective function $Q$, is defined as,
$\mathcal{L}(\mathbf{u},\mathbf{v}^{Q},m,d)=Q(\mathbf{u},m,d)+r(\mathbf{u},\mathbf{v}^{Q},m,d),$ (31)
where $r(\mathbf{u},\mathbf{v},m,d)$ is the weak form in (7). By imposing the condition that the first-order variation of the Lagrangian functional at $\bar{m}$ with respect to the adjoint variable $\mathbf{v}$ and the state variable $\mathbf{u}$ is equal to zero, we obtain, the state problem:
$\textit{Find $\mathbf{u}\in\mathcal{U}$, such that}\quad\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}}\bar{r}\rangle=0,\quad\forall\tilde{\mathbf{v}}\in\mathcal{V},$ (32)
and adjoint problem:
$\textit{Find $\mathbf{v}\in\mathcal{V}$, such that}\quad\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}{\bar{r}^{Q}}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{Q}\rangle,\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$ (33)
In the above relations, $(\bar{\cdot})$ denotes the quantities computed at the mean of the uncertain parameter, and $(\tilde{\cdot})$ indicates directional variables. Directional variables are referred to the variables that define the direction along which perturbation is considered to compute the change in the functional to obtain the derivative of the functional. Also, here we define $r^{Q}=r(\mathbf{u},\mathbf{v}^{Q},m,d)$ for ease of notation. The gradient of $Q$ at $\bar{m}$ acting in the direction $\tilde{m}=m-\bar{m}$ is then given by,
$\langle\tilde{m},\bar{Q}^{m}\rangle=\langle\tilde{m},\partial_{m}\mathcal{L}\rangle.=\langle\tilde{m},\partial_{m}\bar{Q}+\partial_{m}{\bar{r}^{Q}}\rangle\quad\forall\tilde{m}\in\mathcal{M}$ (34)
Thus, evaluating $m$-gradient in (34) requires the solution of state problem (32) for $\mathbf{u}$ and the adjoint problem (33) for $\mathbf{v}$.
To solve for the $m$-Hessian, consider a meta-Lagrangian functional with incremental state $\hat{\mathbf{u}}$ and adjoint $\hat{\mathbf{v}}$
$\small\mathcal{L}^{H}(\mathbf{u},\mathbf{v}^{Q},m,d;\hat{\mathbf{u}}^{Q},\hat{\mathbf{v}}^{Q},\hat{m}^{Q})=\langle\hat{m}^{Q},\partial_{m}\bar{r}^{Q}+\partial_{m}\bar{Q}\rangle+\langle\hat{\mathbf{v}}^{Q},\partial_{\mathbf{v}}\bar{r}^{Q}\rangle+\langle\hat{\mathbf{u}}^{Q},\partial_{\mathbf{u}}\bar{r}^{Q}+\partial_{\mathbf{u}}\bar{Q}\rangle$ (35)
where the terms on the right-hand side are $m$-gradient, forward, and adjoint problems, respectively. Repeating the similar process of vanishing the first variation of the Lagrangian functional with respect to adjoint and state, we obtain the incremental state problem:
$\textit{Find $\hat{\mathbf{u}}\in\mathcal{U}$, such that}\quad\langle\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}\rangle=-\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,\hat{m}^{Q}\rangle,\quad\forall\tilde{\mathbf{v}}\in\mathcal{V},$ (36)
and the incremental adjoint problem:
Find $\hat{\mathbf{v}}\in\mathcal{V}$, such that
$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,\hat{\mathbf{u}}+\partial_{\mathbf{uu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\hat{m}^{Q}+\partial_{\mathbf{u}m}\bar{Q}\,\hat{m}^{Q}\rangle,\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$ (37)
By taking the variation of $\mathcal{L}^{H}$ with respect to $m$ as zero, the Hessian $Q_{mm}(\bar{m})$ acting on $\hat{m}$ is computed as,
$\begin{split}\langle\tilde{m},\bar{Q}^{mm}\,\hat{m}^{Q}\rangle=&\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{Q}\,\hat{\mathbf{v}}^{Q}+\partial_{m\mathbf{u}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{m\mathbf{u}}\bar{Q}\,\hat{\mathbf{u}}^{Q}+\partial_{mm}\bar{r}^{Q}\,\hat{m}^{Q}\\ +\partial_{mm}\bar{Q}\,\hat{m}^{Q}\rangle\end{split}$ (38)
We remark that the last term on the right-hand side of (37) and the last two terms on the right-hand side of (38) are included due to the explicit dependence of the design objective (LABEL:eq:qois) on the uncertain parameter in our problem. Thus, evaluating the $m$-Hessian acting on $\hat{m}$ in (38) requires the solution of the incremental state problem (36) for $\hat{\mathbf{u}}$ and the incremental adjoint problem (37) for $\hat{\mathbf{v}}$.
Similarly, we can compute the gradient and Hessian action of $f$. The adjoint problem can be written as:
$\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{r}^{f}\rangle=-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{f}\rangle,\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}$ (39)
The gradient of $f$ at $\bar{m}$ can be defined as:
$\langle\tilde{m},\bar{f}^{m}\rangle=\langle\tilde{m},\partial_{m}\bar{f}+\partial_{m}\bar{r}^{f}\rangle\quad\forall\tilde{m}\in\mathcal{M}$ (40)
Then by solving incremental state problem,
[…]
and the incremental adjoint problem
[…]
We obtain the Hessian action of $f$ at $\tilde{m}$
$\begin{split}\langle\tilde{m},\bar{f}^{mm}\hat{m}^{f}\rangle=\langle\tilde{m},\partial_{m\mathbf{v}}\bar{r}^{f}\hat{\mathbf{v}}^{f}+\partial_{m\mathbf{u}}\bar{r}^{f}\hat{\mathbf{u}}^{f}+\partial_{m\mathbf{u}}\bar{f}\hat{\mathbf{u}}^{f}+\partial_{mm}\bar{r}^{f}\hat{m}^{f}+\partial_{mm}\bar{f}\hat{m}^{f}\rangle\quad&\\ \forall\tilde{m}\in\mathcal{M}\end{split}$ (43)

### v2

@@ 2.1 Thermomechanical finite element models and design objectives
Figure 2: Finite element solution of the thermomechanical (forward) model for the beam-insulator system with constant porosity $\phi_{f}=0.7$, depicting the solid temperature $\theta_{s}$, fluid temperature $\theta_{f}$, displacement magnitude $||\mathbf{u}_{s}||$, and von Mises stress $T_{\text{VM}}$.
Figure 2 presents the deterministic finite element solution of the forward model for the beam-insulator system under constant porosity $\phi_{f}=0.7$, utilizing the model parameters listed in Table 1. As shown in this figure, thermal expansion of the thermoelastic concrete beam induces stress concentrations within the aerogel thermal break. We define the design objective $Q$ as thermal compliance of the thermal break,
$Q=\frac{1}{2}\sum_{i=s,f}\big{<}\phi_{i}\,\kappa_{i}\,\nabla\theta_{i},\nabla\theta_{i}\big{>}+{\big{<}\phi_{i}h_{\mathrm{conv}}(\theta_{i}-\theta_{\mathrm{a}}),\theta_{i}\big{>}},$ (5)
The design goal is determining the spatial distribution of the fluid volume fraction (porosity), $\phi_{f}(\mathbf{x})$, to minimize stress concentrations within the thermal break, thereby preventing mechanical failure while preserving its thermal insulation performance.
2.2 Spatially correlated uncertain design parameter
Due to the inherent challenges in precisely controlling the properties of additively manufactured aerogel, considerable uncertainty arises in the porosity values within the thermal break arising from the layer-by-layer deposition of aerogel ink. To effectively model such uncertainty, first we define uncertain parameter $m$ as a Gaussian random field characterized by the Matérn covariance kernel. Specifically, $m$ is distributed as $\mathcal{N}(\bar{m},\mathcal{K})$ where $\bar{m}$ is the mean and $\mathcal{K}$ is the covariance operator. The sampling of Gaussian random field $m$ is equivalent to solving the elliptic equation with a robin boundary condition,
@@ 2.2 Spatially correlated uncertain design parameter
Here, $W$ represents the spatial White Gaussian noise with a unit variance, and the parameters $\gamma$ and $\delta$ control the variance $\sigma^{2}={1}/{4\pi\,\delta\,\gamma}$ and spatial correlation length $L_{CR}=\sqrt{{8\gamma}/{\delta}}$. The factor ${\sqrt{\delta\,\gamma}}/{1.42}$ is chosen to minimize boundary artifacts [45], and $\boldsymbol{\Theta}$ is an anisotropic tensor. Then we define the design parameter $d$ as a spatial field, employing a sigmoid mapping function $\phi_{f}\equiv\text{sigmoid}\left(d+m\right)$, ensuring that porosity values asymptotically remain within the physically valid range, maintaining near-linear behavior within $\phi_{f}\in[90\%,10\%]$ [24].
For simplicity, for the rest of this paper, we express the forward state problems (1), (2) and (3) in an abstract form $\mathcal{R}(\mathbf{u},m,d)=0$, and the weak form (4) is expressed as $r(\mathbf{u},\mathbf{v},m,d)$, where $\mathbf{u}$ is the state variables, $\mathbf{v}$ is adjoint variables and $m$ and $d$ are uncertain and design parameters.
@@ 2.3 Chance-constrained PDE-constrained optimal design under uncertainty
We formulate the design problem as minimizing the thermal compliance in (5), which explicitly depends on the uncertain parameter, $Q(d,m)$. The mean-variance cost functional, e.g., [13, 14, 24], is given by:
$\mathcal{J}(d)=\mathbb{E}[Q(d,m)]+\beta_{V}\,\mathbb{V}[Q(d,m)]+\beta_{RG}\,R_{p}(d).$ (7)
where $R_{p}(d)$ is the regularization term, while $\beta_{V}$ and $\beta_{RG}$ represent the coefficients assigned to the variance and the regularization component, respectively.
To prevent mechanical failure of the aerogel thermal break due to stress concentration, the von Mises stress $T_{\text{VM}}$ must remain below a critical threshold. This requirement is incorporated as a probabilistic (chance) constraint in the optimization problem, as $T_{\text{VM}}$ depends on uncertain parameters. However, due to the non-differentiable nature of the maximum stress, direct optimization poses significant challenges. This challenge is addressed using stress aggregation function based on p-norm [33], which is a smooth approximation of the maximum stress and enables gradient-based optimization techniques. The p-norm for the von Mises stress over the domain can be computed as
@@
$f=T_{cr}-T_{\text{pn}}.$ (9)
In above relation $T_{cr}$ is the limiting critical stress, $\mathbf{T}_{0}$ is stress deviator tensor, and $\mathrm{tr}(\cdot)$ is the trace. To avoid stress concentration, we consider a chance constraint:
$P(f(m,d)\geq 0)=\mathbb{E}[\chi_{[0,\infty)}(f(m,d))]\leq\alpha_{c},$ (10)
for a critical chance $10<\alpha_{c}<$ 1, where $\chi_{[0,\infty)}(f(m,d))$ is an indicator function defined as:
$\chi_{[0,\infty)}(f(m,d))=\begin{cases}1&\text{if }f(m,d)\geq 0\\ 0&\text{if }f(m,d)<0\\ \end{cases}$ (11)
Thus, the risk-averse optimal design problem under uncertainty can be formulated as a PDE-constrained optimization problem, expressed as:
$d_{\text{opm}}=\min_{{d}}$ $\ \mathcal{J}({d})$ (12)
@@
$P(f(m,d)\geq 0)\leq\alpha_{c}$
3 Scalable framework for chance constrained optimal design
A conventional method for calculating the mean and variance components in (7) and the probability constrained in (10) is the Monte Carlo estimation for $N_{mc}$ samples,
@@ 3 Scalable framework for chance constrained optimal design
$P(f(.,d)\geq 0)$ $\approx\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\chi_{[0,\infty)]}(f(m^{(i)},d))$ (15)
where $\{m^{(i)}\}_{i=1}^{N_{mc}}$ are random samples from $m$. Achieving an accurate approximation requires a large number of finite element model evaluations due to the slow convergence of the Monte Carlo method, making optimization under uncertainty computationally prohibitive. To address this, we introduce a computationally efficient approach that approximates the mean and variance of the design objective using a quadratic Taylor expansion, leveraging this approximation as a control variate to accelerate Monte Carlo sampling.
3.1 Quadratic Taylor approximation of the design objective and chance constraint
The quadratic Taylor approximation of $Q$ at the mean of the uncertain parameter $\bar{m}$ is given by [24, 13, 14],
$Q_{\text{QUAD}}(m)=\bar{Q}+\big{<}\bar{Q}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{Q}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}.$ (16)
Here, the objective $Q$ and its gradient and Hessian with respect to $m$, evaluated at mean $\bar{m}$, are denoted as $\bar{Q}$, $\bar{Q}^{m}$ and $\bar{Q}^{mm}$, respectively. The expectation and variance of $Q_{\text{QUAD}}$ are derived in [24] such that,
$\mathbb{E}[Q_{\text{QUAD}}]=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{k}\big{)},\quad\mathbb{V}[Q_{\text{QUAD}}(m)]=\big{<}\bar{Q}^{m},\mathcal{K}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{k})^{2}),$ (17)
where ¯$\bar{\mathcal{H}}_{k}=\mathcal{K}\,\bar{Q}^{mm}$ represents the preconditioned Hessian for covariance at the mean of the uncertain parameter $m$ and $\mathcal{K}$ is the covariance operator. The design objective in (7) can then be approximated as
$\mathcal{J}_{\mathrm{QUAD}}(d)=\bar{Q}+\frac{1}{2}\mathrm{tr}\big{(}\bar{\mathcal{H}}_{k}\big{)}+\beta_{V}\,\bigg{(}\big{<}\bar{Q}^{m},\mathcal{K}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\mathrm{tr}((\bar{\mathcal{H}}_{k})^{2})\bigg{)}+\beta_{RG}\,R_{p}(d).$ (18)
@@ 3.1 Quadratic Taylor approximation of the design objective and chance constraint
We can also construct a second-order Taylor expansion of the chance constraint function $f$ at $\bar{m}$ as,
$f_{QUAD}(m)=\bar{f}+\big{<}\bar{f}^{m},\,m-\bar{m}\big{>}+\frac{1}{2}\big{<}\bar{f}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>},$ (19)
@@
where $\bar{f}$, $\bar{f}^{m}$ and $\bar{f}^{mm}$ are the constraint function $f$ and it gradient and Hessian evaluated at $\bar{m}$. The probability $P(f(.,d)\geq 0)$ by Monte Carlo estimator (15) with the quadratic approximation yields,
$P(f(m,d))\geq 0)=\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\chi_{[0,\infty)}(f_{\text{QUAD}}(m_{i},d)).$ (20)
In order to compute the second-order Taylor expansion (19), we need to compute the term $\big{<}\bar{f}^{mm}\,(m-\bar{m}),\,m-\bar{m}\big{>}$, which can be approximated through a low-rank approximation as,
$\big{<}\bar{f}^{mm}(m-\bar{m}),m-\bar{m}\big{>}=\sum_{i=1}^{N_{\mathrm{eig}}^{f}}\lambda_{i}^{f}\big{|}\big{<}m-\bar{m},\mathcal{K}^{-1}\psi_{i}^{f}\big{>}\big{|}^{2},$ (21)
where $\lambda_{i}^{f}$ and $\psi_{i}^{f}$ are the eigenvalues and eigenfunctions corresponding to the solution of a generalized eigenvalue problem. The generalized eigenvalue problem corresponding to the second-order Taylor expansion can be formulated as,
$\mathcal{K}^{1/2}\>\bar{f}^{mm}\>\mathcal{K}^{1/2}\>\psi_{i}^{f}=\lambda_{i}^{f}\psi_{i}^{f},\quad i=1,2,\dots,N_{\mathrm{eig}}^{f},$ (22)
where $\lambda_{1}^{f},\lambda_{2}^{f},...,\lambda_{N_{\mathrm{eig}}^{f}}^{f}$ are the $N_{\mathrm{eig}}^{f}$ largest eigenvalues. Computing the approximations of the mean, variance, and constraint function requires a Lagrangian-based formalism to derive the $m$-gradient, $m^{f}$-gradient, $m$-Hessian and $m^{f}$-Hessian, which represent the gradients and Hessians of the design objective $Q$ and constraint function $f$ with respect to $m$, as detailed in Section A.
Finally, the the mean and variance in (17) requires the trace of the covariance-preconditioned Hessian ¯$\bar{\mathcal{H}}_{k}$. Here, we leverage a randomized trace estimator using the dominant eigenvalues $\{\lambda^{Q}_{n}\}_{n=1}^{N^{Q}_{\mathrm{eig}}}$ of ¯$\bar{\mathcal{H}}_{k}$,
$\mathrm{tr}\big{(}\bar{\mathcal{H}}_{k}\big{)}\approx\sum_{n\geq 1}^{N^{Q}_{\mathrm{eig}}}\lambda^{Q}_{n},\quad\mathrm{tr}\big{(}\bar{\mathcal{H}}_{k}^{2}\big{)}\approx\sum_{n\geq 1}^{N^{Q}_{\mathrm{eig}}}{(\lambda^{Q}_{n})}^{2}.$ (23)
When eigenvalues decay rapidly, $N^{Q}_{\mathrm{eig}}$ only depend on the intrinsic low dimensionality of the otherwise high-dimensional parameters. The eigenvalues $\lambda^{Q}_{n}$ are obtained by solving the generalized eigenvalue problem with an input of oversampling factor $N_{o}$, via double-pass randomized eigenvalue solver (see [24, 14] for more details).